. A body executing S.H.M. along a straight line has a velocity of 3 ms–1 when it is at a distance of 4 m from its mean position and 4 ms–1 when it is at a distance of 3 m from its mean position. Its angular frequency and amplitude are

Grade:12th pass

1 Answers

venkat
5 years ago
v=\omega \sqrt{a^{2}-x^{2}}
v_{1}=\omega \sqrt{a^{2}-x_{1}^{2}}
3=\omega \sqrt{a^{2}-4^{2}}
9=\omega ^{2}(a^{2}-16)
\omega ^{2}=9/(a^{2}-16)​                -eqn(1)
v_{2}=\omega \sqrt{a^{2}-x_{2}^{2}}
4=\omega \sqrt{a^{2}-{3}^{2}}
16=\omega^{2} ({a^{2}-9})​                   -eqn-(2)
After simplifyng  eqns-(1) & (2) you will get, 
Angular frequency \ \omega = 1 \ rad/s
Amplitude \ a \ = \ 5m
Approve my answer if it helped you.